Find the solution of the initial value problem y′′+7y′+10y=0, y(0)=11 and y′(0)=−46.
Answer #1

Solution or y+ 7y +10 =0 -, bet y=mx be the solution of ta from ③ we have (m² + 7m +10) mxo 6 MX m2 + 7m + 10 =0 m² + sm +.. १५5 = 46 १८+ १ २ = १५ 3 2 = 24 (28) C+ 8 = 11 (C-30 -2 -Sx Y(M) = 3e +8e)